For the sublimation of iodine at 25C

I2(s) yield I2(g)

the values of H and G are, respectively, 62 kJ and 19 kJ. Estimate the temperature at which iodine sublimes. Assume H and S do not depend on temperature.
Answer #1

Given,

For the sublimation reaction of iodine at 25°C

standard enthalpy, ΔH = 62 kJ = 62000 J

standard free energy ΔG = 19 kJ = 19000 J

T = 273.15 + 25 = 298.15 K

Using these values and below formula, calculate the standard entropy of 25°C

ΔG = ΔH-TΔS

19000 J = 62000 J - (298.15 K)ΔS

ΔS = (62000 -19000)J/(298.15)K

ΔS = 144.22 J/K

Calculate the temperature for the sublimation of iodine using following formula

T = ΔH/ΔS

T = (62000 J)/(144.22 J/K)

T = 429.9 K

T = 429.9-273.15 = 156.8°C

Temperature for sublimation of iodine = 156.8°C or 429.9 K